i actually have two questions.
1) what is the answer to this question: the sum of two consecutive is -67
find the integers
2) the sum of two consecutive integers is 28
find the integers
how do i do this and write it out?

Answer by Alan3354(21612)

1) what is the answer to this question: the sum of two consecutive is -67
find the integers
They have to be approximately -67/2. -67/2 = -33.5, so they're -33 and -34
------------------
The hard way:
n + (n+1) = -67
2n+1 = -67
2n = -68
n = -34
--> -34 and -33
------------------
2) the sum of two consecutive integers is 28
find the integers
It's not possible for the sum of 2 consecutive integers to be an even number.
-----------------------------
how do i do this and write it out?